Q-1: Given the following code fragment, what is its Big-O running time?

test = 0

for i in range(n):

            for j in range(n):

                        test= test + i *j

Q-2: Given3 the following code fragment what is its Big-O running time?

for i in range(n):

            test=test+1

for j in range(n):

            test= test - 2

Q-3: Given the following code fragment what is its Big-O running time?

i = n

while i > 0:

            k=2+2

            i = i // 2
Answer #1

Here is the answer..

testo for ein range (s): for j in range(n) - test = lest + 1 tj nxn nxo 1+0+0+000 3 if (1) for &(n)= ant+n+1 flu)= 0 (n) a in

If you have any doubts please COMMENT...

If you understand the answer please give THUMBS UP...